The denver post stated that 80% of all new products introduced in grocery stores fail within in 2 years.If a grocery store chain introduces 66 new products,what is the probability that within 2 years 47 or more will fail?
Answer #1

The denver post stated that 80% of all new products introduced in grocery stores fail within in 2 years.
probability of failure, p = 0.8
If a grocery store chain introduces 66 new products, sample size, n = 66

mean no. of products introduced in grocery stores that fail within in 2 years,

μ = n*p = 66*0.8 = 52.8

probability that within 2 years 47 or more will fail, P(x47)

=P()

=P(z -0.09)

=1-0.4641=0.5359

answered by: madzia